Question 47
ProbabilityIn an athletics competition, the probability that an athlete wins a 100 m race is 1/8 and the probability that he wins in high jump is 1/4. What is the probability that he wins only one of the events?
- A.3/32
- B.3/16
- C.7/32
- D.5/16
Show the answer and explanation
Answer: D (5/16)
Explanation
P(race only) = 1/8 × 3/4 = 3/32; P(jump only) = 7/8 × 1/4 = 7/32. Total = 10/32 = 5/16.